# SafeNlidb: A Privacy-Preserving Safety Alignment Framework for LLM-based Natural Language Database Interfaces

### v1

@@ Introduction
    Safe SQL Set Synthesis. This includes two subtypes, soft safe is created by modifying unsafe SQL queries through counterfactual reasoning (e.g., replacing critical SQL queries to invalidate the overall attack), while hard safe is extracted and composed from existing harmless datasets.
@@
Reasoning Warm-Up and Alternating Preference Optimization (APO)
@@ A.3 Preference Data Collection
Previous work (liu2025uncoveringimpactchainofthoughtreasoning) has pointed out the importance of constructing preference data containing fine-grained reasoning processes to improve the Text-to-SQL capabilities of models, and they inspire our preference data collection process. Specifically, for a model after the reasoning warm-up, we let it sample $N$ candidate answers for each sample in the training set. Then, according to the correctness of security and SQL execution results in these candidate answers, they are divided into preference data containing chosen samples and rejected samples (§ Reasoning Warm-Up and Alternating Preference Optimization (APO)).
@@ A.5 Reliability Score (RS) Evaluation
$\phi(x)=\begin{cases}\,\,1&\text{if }s(x){=}1,{g}(x){=}1,\text{EX}(x){=}1,\\ {-}0.5&\text{if }s(x){=}1,{g}(x){=}1,\text{EX}(x){=}0,\\ {-}1&\text{if }s(x){=}1,{g}(x){=}0,\\ {-}1&\text{if }s(x){=}0,{g}(x){=}0\end{cases}$ (4)
Here, $s(x)$ denotes the security label of $x$, where 1 indicates safe and 0 indicates unsafe. The function $g(x)$ evaluates the model’s prediction correctness regarding the security of $x$, with 1 representing a correct prediction and 0 representing an incorrect prediction. $\text{EX}(x)$ measures the execution accuracy of the generated SQL (1 if the execution results match the ground truth, 0 otherwise) (zhong-etal-2020-semantic). After computing sample-level scores, the final Reliability Score is defined as $\text{RS}=\frac{1}{N_{T}}\sum_{x}\phi(x)$, where $N_{T}$ represents the total number of samples.
@@ C.2 Post-Hoc Detection
Method TS
Further Pre-Training Methods
Codes-1B (DS: 21.5GB) (10.1145/3654930) 55.7
Codes-3B (DS: 21.5GB) (10.1145/3654930) 64.6
Codes-7B (DS: 21.5GB) (10.1145/3654930) 66.0
Codes-15B (DS: 21.5GB) (10.1145/3654930) 69.3
OmniSQL-7B (DC: 2.5M) (li2025omnisqlsynthesizinghighqualitytexttosql) 76.9
Prompt-Based Methods
GPT-4 (Few-shot) (DIN-SQL) 67.4
C3 + ChatGPT (dong2023c3zeroshottexttosqlchatgpt) 71.4
DIN-SQL + GPT-4 (DIN-SQL) 74.2
DAIL-SQL + GPT-4 (dail-sql) 76.6
SQL-PaLM (Few-shot) (sun2024sqlpalm) 77.3
DAIL-SQL + GPT-4 + SC (dail-sql) 76.2
GPT-4o-mini (li2025omnisqlsynthesizinghighqualitytexttosql) 70.4
GPT-4-Turbo (li2025omnisqlsynthesizinghighqualitytexttosql) 72.4
GPT-4o (li2025omnisqlsynthesizinghighqualitytexttosql) 70.9
Llama3-8B 60.7
SQL-CoT + Llama3-8B (DS: <<500MB, DC: 28K) 70.0
Qwen2.5-7B 73.1
SQL-CoT + Qwen2.5-7B (DS: <<500MB, DC: 28K) 76.3
Table 6: Comparison results of our method with the further specific pre-training methods and closed-source LLM-based methods on interactive scenarios constructed from the Spider-Dev dataset. DS: Sample storage size. DC: Sample set size. SC: Self-Consistency strategy.
Method EX
Further Pre-Training Methods
Codes-1B (DS: 21.5GB) (10.1145/3654930) 22.2
Codes-3B (DS: 21.5GB) (10.1145/3654930) 29.1
Codes-7B (DS: 21.5GB) (10.1145/3654930) 30.8
Codes-15B (DS: 21.5GB) (10.1145/3654930) 34.1
Prompt-Based Methods
Codex 175B (bird) 25.4
ChatGPT (bird) 24.1
ChatGPT + CoT (bird) 25.9
Claude-2 (bird) 28.3
GPT-4 (bird) 30.9
Llama3-8B 19.9
SQL-CoT + Llama3-8B (DS: <<500MB, DC: 28K) 23.0
Qwen2.5-7B 23.3
SQL-CoT + Qwen2.5-7B (DS: <<500MB, DC: 28K) 26.0
Table 7: Comparison of our method with further specific pre-training methods and closed-source LLM-based methods on interactive scenarios constructed from the Bird-Dev (without external knowledge) dataset. DS: Sample storage size. DC: Sample set size.
@@ D More Results and Analysis
To further verify the effectiveness of the SQL-CoT proposed in our framework, we compare the performance of different methods on two authoritative Text-to-SQL benchmarks: Spider (yu-etal-2018-spider) and Bird (bird). It is worth emphasizing that SQL‑CoT does not rely on any training data officially provided by these benchmarks. As shown in Table 6 and Table 7, SQL-CoT significantly improves model performance on both general text-to-SQL tasks, achieving performance gains of 9.3%/3.1% on Llama3-8B and 3.2%/2.7% on Qwen2.5-7B, respectively. Notably, our approach can even achieve leapfrogging: its performance is better than larger, further pre-trained, and even partially closed-source LLMs. SQL‑CoT + Llama3‑8B outperforms the Codes‑15B model (70.0 vs. 69.3), while SQL‑CoT + Qwen2.5‑7B delivers results comparable to or even better than methods based on GPT‑4, such as DIN‑SQL and DAIL‑SQL. Furthermore, Table 9 presents a comprehensive performance comparison of various models. The results demonstrate that SQL‑CoT effectively enhances the capabilities of smaller-scale LLMs, significantly narrowing the performance gap with their larger-scale counterparts. In certain scenarios, it even enables smaller models to surpass much larger ones. A notable example is that SQL‑CoT combined with Qwen2.5‑7B achieves higher performance than both Qwen2.5‑14B and Qwen2.5‑72B on the Spider dataset. The above results highlight the strong generalization capability of our synthetic data methodology, which not only reduces data and computational costs substantially but also extends model capabilities beyond conventional Text-to-SQL to include enhanced protection of sensitive data. Furthermore, these results provide additional evidence for the crucial role of properly designed CoT in strengthening LLMs’ complex reasoning abilities (liu-etal-2023-uncovering; li2025omnisqlsynthesizinghighqualitytexttosql).
Model SecureSQL ShieldSQL
Llama3-8B
SQL-CoT 58.75 67.74
Safenlidb 61.80 66.67
$\Delta$ (+3.05) (-1.07)
Qwen2.5-7B
SQL-CoT 60.32 66.94
Safenlidb 58.59 65.00
$\Delta$ (-1.73) (-1.94)
@@
In previous experiments, we demonstrated the generalization ability of the synthesized SQL-CoT for the Text-to-SQL task. Building on that, this analysis examines the impact of incorporating Safety-CoT into SQL-CoT to form H-CoT on Text-to-SQL performance. As shown in the results from Table 8, while H-CoT introduces additional safety considerations, its performance remains comparable to SQL-CoT, with only a slight decline of about 1%. Notably, when using Llama3-8B as the baseline, H-CoT shows a +3.05% improvement over SQL-CoT on SecureSQL-SQL. These results confirm that our approach effectively integrates safety evaluation capabilities into models while maintaining their core SQL generation proficiency.
@@
Dataset Name Avg. Rounds #Unsafe Type #DB Constraint Types #Total
Column/Other Num/Str
SecureSQL 1.15 3 57 932
ShieldSQL 1.95 9 455 237/303 270/270 540
Synthetic Data 1.66 9 2,000 12,276/16,607 14,981/13,902 28,883
Table 11: Quantitative statistics of different datasets. Avg. Rounds indicates the average number of interaction turns, #Unsafe Type enumerates the count of insecure interaction patterns, and #DB specifies the number of databases. Constraint Types categorizes security constraints through two dimensions: (1) by scope (column-level vs. other types including row-level or hybrid row-column constraints) and (2) by value type (numeric vs. string). #Total reports the total sample size for each dataset.

### v2

@@ Introduction
    Benign SQL Set Synthesis. This includes two subtypes, soft safe is created by modifying unsafe SQL queries through counterfactual reasoning (e.g., replacing critical SQL queries to invalidate the overall attack), while hard safe is extracted and composed from existing harmless datasets.
@@
Alternating Preference Optimization (APO)
@@ A.2 Interaction Sample Synthesis
Algorithm 1 Alternating Preference Optimization (APO)
0: Dataset $\mathcal{D}_{\text{sft}}$; warmed-up model $\pi_{\text{SFT}}$; safety verifier $\mathcal{V}_{\text{safe}}$; SQL execution verifier $\mathcal{V}_{\text{exec}}$; rollout size NN
1:for each $x\in\mathcal{D}_{\text{sft}}$ do
2:  Generate candidate responses $\{y_{i}\}_{i=1}^{N}\sim\pi_{\text{SFT}}(\cdot\mid x)$
3:  Evaluate safety and execution of each candidate
4:  Chosen set:
$\mathcal{Y}_{w}\leftarrow\{\,y_{i}:\mathcal{V}_{\text{safe}}(y_{i})=1\land\mathcal{V}_{\text{exec}}(y_{i})=1\,\}$
5:  Rejected set: $\mathcal{Y}_{l}\leftarrow\{\,y_{i}:y_{i}\notin\mathcal{Y}_{w}\,\}$
6:  Priority strategy for rejected samples: ➀Select candidates with both safety and SQL incorrect:
$\mathcal{Y}_{l}^{\text{both}}\leftarrow\{y_{i}\in\mathcal{Y}_{l}:\mathcal{V}_{\text{safe}}(y_{i})=0\land\mathcal{V}_{\text{exec}}(y_{i})=0\}$
➁ If empty, select candidates with a single error:
$\mathcal{Y}_{l}^{\text{single}}\leftarrow\{y_{i}\in\mathcal{Y}_{l}:\mathcal{V}_{\text{safe}}(y_{i})=0\lor\mathcal{V}_{\text{exec}}(y_{i})=0\}$
➂If both are empty, skip this sample
7:  For $y_{l}$ from $\mathcal{Y}_{l}^{\text{single}}$, anchor correct reasoning spans using the corresponding segments from ${y}_{w}$
8:  Get preference pairs $(x,y_{w},y_{l})$ with $y_{w}\in\mathcal{Y}_{w}$, $y_{l}\in\mathcal{Y}_{l}$
9:end for
10: Update $\pi_{\text{APO}}$ via the objective function $\mathcal{L}_{\text{APO}}$
@@ A.3 Preference Data Collection
Previous work (liu2025uncoveringimpactchainofthoughtreasoning) has pointed out the importance of constructing preference data containing fine-grained reasoning processes to improve the Text-to-SQL capabilities of models, and they inspire our preference data collection process. Specifically, for a model after the reasoning warm-up, we let it sample $N$ candidate answers for each sample in the training set. Then, according to the correctness of security and SQL execution results in these candidate answers, they are divided into preference data containing chosen samples and rejected samples (§ Alternating Preference Optimization (APO)). The detailed algorithmic process of APO is illustrated in Algorithm 1.
@@ A.5 Reliability Score (RS) Evaluation
$\phi(x)=\begin{cases}\,\,k_{1}&\text{if }s(x){=}1,{g}(x){=}1,\text{EX}(x){=}1,\\ \,\,k_{2}&\text{if }s(x){=}1,{g}(x){=}1,\text{EX}(x){=}0,\\ \,\,k_{3}&\text{if }s(x){=}1,{g}(x){=}0,\\ \,\,k_{4}&\text{if }s(x){=}0,{g}(x){=}0\end{cases}$ (4)
Here, following (lee2024trustsql), we set $k_{1}=k_{3}=k_{4}=-1$ and $k_{2}=-0.5$. $s(x)$ denotes the security label of $x$ (1 for safe, 0 for unsafe). The function $g(x)$ evaluates the model’s prediction correctness regarding the security of $x$ (1 if correct, 0 if not). $\text{EX}(x)$ measures SQL execution accuracy (1 if results match the ground truth, 0 otherwise) (zhong-etal-2020-semantic). The final score is defined as $\mathrm{R}=\frac{1}{N_{T}}\sum_{x}\phi(x)$, $N_{T}$ is the total number of samples.
@@ C.2 Post-Hoc Detection
Method DS Spider Bird
Single-Turn: Closed-Source LLMs (Prompt-Based)
GPT-3.5-turbo (bird) 28.0
GPT-3.5-turbo + C3 (dong2023c3zeroshottexttosqlchatgpt; 10.1145/3654930) 71.4
GPT-4 (DIN-SQL; bird) 67.4 30.9
GPT-4 + DIN-SQL (DIN-SQL; 10.1145/3654930) 74.2
GPT-4 + DAIL-SQL (dail-sql; 10.1145/3654930) 76.6
SQL-PaLM (Few-shot) (sun2024sqlpalm; 10.1145/3654930) 77.3
GPT-4 + DAIL-SQL + SC (dail-sql; 10.1145/3654930) 76.2
GPT-4o-mini (li2025omnisqlsynthesizinghighqualitytexttosql) 70.4 58.8
GPT-4-Turbo (li2025omnisqlsynthesizinghighqualitytexttosql) 72.4 62.0
GPT-4o (li2025omnisqlsynthesizinghighqualitytexttosql) 70.9 61.9
Single-Turn: Open-Source LLMs (Pre-Trained-Based)
Codes-1B (10.1145/3654930) 21.5 55.7 22.2
Codes-3B (10.1145/3654930) 21.5 64.6 29.1
Codes-7B (10.1145/3654930) 21.5 66.0 30.8
Codes-15B (10.1145/3654930) 21.5 69.3 34.1
OmniSQL-7B (li2025omnisqlsynthesizinghighqualitytexttosql) 9.4 76.9 55.1
Multi-Turn: Open-Source LLMs
Llama3-8B - 60.7 19.9
SQL-CoT + Llama3-8B 0.2 70.0 23.0
Qwen2.5-7B - 73.1 23.3
SQL-CoT + Qwen2.5-7B 0.2 76.3 26.0
Table 6: Comparison of different methods on the general Text-to-SQL tasks. DS (GB) denotes the training sample storage size. The Spider and Bird datasets use dev sets, with TS (zhong-etal-2020-semantic) and EX (bird) as evaluation metrics, respectively. $\dagger$ represents results in multi-turn settings (which involve more complex context compared to single-turn interactions).
Method Security SecureSQL-$\mathrm{EX}$ ShieldSQL-$\mathrm{EX}$
Llama3-8B
SQL-CoT
59.80 41.35
SafeNlidb
62.75 (+2.95+2.95) 46.62 (+5.27+5.27)
Qwen2.5-7B
SQL-CoT
56.87 44.36
SafeNlidb
59.48 (+2.61+2.61) 48.12 (+3.76+3.76)
@@ D More Results and Analysis
To further verify the effectiveness of the SQL-CoT proposed in our framework, we compare the performance of different methods on two authoritative Text-to-SQL benchmarks: Spider (yu-etal-2018-spider) and Bird (bird). It is worth emphasizing that SQL‑CoT does not rely on any training data officially provided by these benchmarks. As shown in Table 6, SQL-CoT significantly improves model performance on both general text-to-SQL tasks, achieving performance gains of 9.3%/3.1% on Llama3-8B and 3.2%/2.7% on Qwen2.5-7B, respectively. Notably, our approach can even achieve leapfrogging: its performance is better than larger, further pre-trained, and even partially closed-source LLMs. SQL‑CoT + Llama3‑8B outperforms the Codes‑15B model (70.0 vs. 69.3), while SQL‑CoT + Qwen2.5‑7B delivers results comparable to or even better than methods based on GPT‑4, such as DIN‑SQL and DAIL‑SQL. Furthermore, Table 9 presents a comprehensive performance comparison of various models. The results demonstrate that SQL‑CoT effectively enhances the capabilities of smaller-scale LLMs, significantly narrowing the performance gap with their larger-scale counterparts. In certain scenarios, it even enables smaller models to surpass much larger ones. A notable example is that SQL‑CoT combined with Qwen2.5‑7B achieves higher performance than both Qwen2.5‑14B and Qwen2.5‑72B on the Spider dataset. The above results highlight the strong generalization capability of our synthetic data methodology, which not only reduces data and computational costs substantially but also extends model capabilities beyond conventional Text-to-SQL to include enhanced protection of sensitive data. Furthermore, these results provide additional evidence for the crucial role of properly designed CoT in strengthening LLMs’ complex reasoning abilities (liu-etal-2023-uncovering; li2025omnisqlsynthesizinghighqualitytexttosql).
@@
In the previous section, we demonstrated the generalization ability of the synthesized SQL-CoT for the Text-to-SQL task. In this section, we will further investigates how introducing safety reasoning affects SafeNlidb’s core Text-to-SQL performance. As shown in the results from Table 7, although SafeNlidb introduces an additional safety reasoning process, it still consistently outperforms the baseline that uses only SQL-CoT, achieving 2.61% to 5.27% gains in Text-to-SQL performance across both datasets. This further proves that our SafeNlidb framework seamlessly integrates safety reasoning capabilities into the Text-to-SQL model, enhancing core SQL generation capability while ensuring multi-turn interaction safety.
Figure 8: Comparison of Reliability Scores ($\mathrm{R}$) between SafeNlidb and DP methods (blue) under different reward and penalty parameter settings ($k_{1}$ and $k_{2}$). $\uparrow$ represents the spatial upward or downward relation of the DP method’s result plane relative to that of ours. The gradient-colored areas indicate where our method outperforms the DP method.
Method SecureSQL ShieldSQL
[…]
$\mathrm{S}$ $\mathrm{EX_{SAFE}}$ $\mathrm{R}$ $\mathrm{0.1}$ $\mathrm{0.3}$ $\mathrm{0.5}$ $\mathrm{0.7}$ $\mathrm{0.9}$ $\mathrm{1.0}$ $\mathrm{S}$ $\mathrm{EX_{SAFE}}$ $\mathrm{R}$ $\mathrm{0.1}$ $\mathrm{0.3}$ $\mathrm{0.5}$ $\mathrm{0.7}$ $\mathrm{0.9}$ $\mathrm{1.0}$
Differential Privacy Upper Bound (with Ground-Truth SQL)
DP $\sim{100}$ 0 -16.4 10.08 3.25 2.01 1.41 1.12 1.04 $\sim{100}$ 0 -15.8 9.94 3.13 2.10 1.36 1.03 1.03
Differential Privacy (with LLMs)
$\text{DP}_{\text{Llama3-8B}}$ $\sim\textbf{{100}}$ 0 -16.4 16.18 5.40 3.25 2.30 1.80 1.61 $\sim\textbf{100}$ 0 -15.8 20.01 6.98 3.88 2.79 3.24 1.95
$\text{DP}_{\text{Qwen2.5-7B}}$ $\sim\textbf{{100}}$ 0 -16.4 14.76 4.96 2.97 2.13 1.65 1.46 $\sim\textbf{100}$ 0 -15.8 22.55 7.52 4.47 3.27 2.49 2.26
Backbone LLMs
Llama3-8B 54.9 29.41 -40.7 18.61 6.23 3.74 2.67 2.08 1.86 52.4 26.90 -43.9 14.61 4.87 2.92 2.08 1.62 1.46
Qwen2.5-7B 52.2 57.19 -35.0 11.52 3.84 2.30 1.65 1.28 1.15 69.1 18.13 -36.9 19.80 6.61 3.96 2.83 2.20 1.98
Llama3-70B 58.4 31.05 -35.0 11.08 3.70 2.21 1.58 1.23 1.11 64.8 33.92 -43.7 10.95 3.64 2.19 1.56 1.21 1.10
SafeNlidb
$\textsc{Ours}_{\text{Llama3-8B}}$ 64.4 50.98 -24.4 10.43 3.47 2.08 1.49 1.15 1.04 84.6 33.92 -13.8 9.57 3.19 1.91 1.36 1.06 0.95
$\textsc{Ours}_{\text{Qwen2.5-7B}}$ 63.1 48.37 -28.2 10.86 3.62 2.17 1.55 1.21 1.08 81.9 35.67 -15.8 8.76 2.92 1.75 1.25 0.97 0.88
Table 8: Comprehensive comparison between the DP method from both the Text-to-SQL and database perspectives. $\diamondsuit$ represents using Ground-Truth SQL directly. $\textit{EX}_{\textit{SAFE}}$} denotes the execution accuracy on benign samples. $\sim{{100}}$ indicates the ideal security upper bound of the DP method. The best result is highlighted in bold, and the second-best result is underlined.
Comparison Analysis with Differential Privacy.
In the previous analyses, we have demonstrated that the internal alignment approach SafeNlidb outperforms rule-based database defense strategies (e.g., SQD, SSA, DEM). In this section, we further introduce Differential Privacy (DP), a representative noise-based database defense method, and provides a comprehensive comparison from both multi-turn Text-to-SQL and database SQL execution perspectives.
Multi-turn Text-to-SQL Security and Reliability. In addition to Security Accuracy and Reliability Score used in this work, we introduce $\mathrm{EX_{SAFE}}=0$ to evaluate the execution accuracy of benign queries. As shown in Table 8, although DP theoretically achieves a higher security upper bound and obtains a high score in Reliability Score (R), its noise injection mechanism in SQL query results significantly reduces execution accuracy ($\mathrm{EX_{SAFE}}=0$). In Text-to-SQL tasks where execution accuracy is crucial, such a trade-off between precision and safety greatly limits its practicality. In contrast, SafeNlidb, through internal safety alignment within the LLM, enhances security while maintaining high SQL execution accuracy. Particularly on the ShieldSQL dataset, it achieves reliability scores comparable to or even exceeding those of DP. We also note that although Qwen2.5-7B performs close to random in terms of security, its raw Text-to-SQL performance is quite good ($\mathrm{EX_{SAFE}}=57.19$). We speculate that this capability may stem from its extensive pre-training on code-related tasks (qwen2025qwen25technicalreport). Overall, our method successfully achieves a balance between security and usability.
Figure 8 shows the comparison of Reliability Scores between our method and the DP method under different reward and penalty parameters ($k_{1}$ and $k_{2}$), which are the only parameters affecting the DP method. The result plane of the two methods intersect in 3D space. Overall, SafeNlidb demonstrate a clear overall advantage over the DP method. On the SecureSQL dataset, our method not only performs similarly to DP, but in some cases, it even surpasses DP. On the ShieldSQL dataset, this advantage is further expanded, outperforming the DP method in the majority of scenarios. This further highlights the potential of SafeNlidb in enhancing the security and reliability of Text-to-SQL tasks.
SQL Execution Mean Absolute Error. From the database perspective, we further compare both approaches using the Mean Absolute Error (MAE) (10.14778/3594512.3594515), a widely used metric in differential privacy, to measure the average deviation between noisy and true query results. This metric directly reflects the loss in query utility introduced by differential privacy noise. The calculation is as follows:
$\mathrm{MAE}=\frac{1}{N}\sum_{i=1}^{N}\left|\tilde{{r}}_{i}-{{r}}_{i}\right|$ (5)
Where $\tilde{{r}}_{i}$ denotes the noisy SQL query result and ${{r}}_{i}$ represents the true result. For fair comparison, erroneous SQL generated by the LLM is assigned the maximum noise based on the true execution result. Following prior work (10.14778/3594512.3594515), each SQL is perturbed 100 times, and the average result is reported.
Results in Table 8 show that SafeNlidb consistently outperforms both pure LLMs and LLM+DP combinations in terms of MAE across all privacy budgets. Notably, our method achieves MAE performance comparable to or even surpassing the DP upper bound (using Ground-Truth SQL). These results strongly validate the comprehensive advantage of our method over pure database defense strategies in multi-turn Text-to-SQL scenarios.
@@
Dataset Rationale Split Avg.Rounds #Unsafe Type #DB Constraint Types #Total
(CoT) Column/Others Num/Str
SecureSQL
Test 1.15 3 57 932
ShieldSQL
Test 1.95 9 455 237/303 270/270 540
Synthetic Data
Train 1.66 9 2,000 12,276/16,607 14,981/13,902 28,883
Table 11: Quantitative statistics of different datasets. Avg.Rounds indicates the average number of interaction turns. #Unsafe Type enumerates the count of insecure interaction patterns, and #DB specifies the number of databases. Constraint Types categorizes security constraints through two dimensions: (1) scope (column-level vs. other row-level or hybrid row-column constraints) and (2) value type (numeric vs. string). #Total reports the total number of samples in each dataset.